OpenStudy (aizhalee):

Can someone help me please , Medal and Fan will be given what is the point that divides the segment a(-10,-2) and b(7,8) into the ratio 2:3

OpenStudy (aizhalee):

Thank you Niels

OpenStudy (anonymous):

you are trying to find a point 2/5 of the way from a to b, 7--10=17 =the distance from a to b 17/5*2=6.8 = 2/5 into that distance 6.8+-10=-3.2 =2/5 of the distance + the original amount 8--2=10 10/5*2=4 4+-2=2 the point is (-3.2,2) do you understand it like this?

OpenStudy (aizhalee):

so that proves the ratio is2:3 ?

OpenStudy (aizhalee):

Thank you for explaining I appreciate it

OpenStudy (anonymous):

no problem! the distance from A to this point devided by the distance from this point to B is 2/3

OpenStudy (aizhalee):

Thank you God bless :) Youre a life saver !! you have no gotten a new fan :D

OpenStudy (anonymous):

haha thank you!